OpenStudy (sighn0more):

Hello! Can someone please help me work through this quadratic equation which must be solved through completing the square?: x^2+3+4 = 0

OpenStudy (sighn0more):

So far I have moved the constant to the RHS x^2 + 3 + ___ = -4 + ___

OpenStudy (sighn0more):

What I'm confused on is the number which will fill the blank space. It should be 2.25 but this would make factoring difficult.

OpenStudy (tkhunny):

Think hard on this: \((a+b)^{2} = a^{2} + 2ab + b^{2}\) Don't use decimals. (3/2)^2 = 9/4 It will NEVER make factoring difficult. You are Completing the square. It is, by design, easy to factor after that. The factoing will always be \((a+b)^{2}\) - every time.

OpenStudy (sighn0more):

Oh, I see now. So it will be (x + 3)^2 and then I will add 9/4 to 4. Thank you!!